6. For the reaction 2 HBr(g) Cl2(g) 2 HCl(g) Brz(g) a. Write the equilibrium constant expression for the reaction. b. Using the following G values, calculate AG for the reaction. HCl(g) -95.27 mol HBr(g) 53.22 mol Bra(g) 3.14 Cl2(g)* 0 mol c. Calculate the equilibrium constant Keq be at 298 K. d. Does this equilibrium lie more with reactants or products? udant u have neither given nor received
Answer #1

a)

K = [products]^p /[ reactants]^r

K = [HCl]^2[Br] /([HBr]^2[Cl2]

b)

dG = Gproducts - Greactants

dG = 2*-95.27 + 3.14 - (2*-53.22 + 0)

dG = -80.96 kJ/mol

c)

Keq --> dG = -RT*ln(Keq)

Keq = exp(-dG/(RT))

Keq = exp(80960/(8.314*298))

Keq = 1.55420*10^14

d)

this favours strongly the products